OpenStudy (volleyballlover55):

please help fast! If a spacecraft was parked on Venus and needed to make a flight to Jupiter, how far would it need to travel? Show your work and provide your answer in scientific notation.

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (mayankdevnani):

just find distance between them

OpenStudy (countryboy15):

The distance varies between 670.1 million kilometers at their closest and 1172.7 million kilometers at their farthest. hope this helps also please medal if it did thanks have a great day

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (mayankdevnani):

just google the distance between them

OpenStudy (volleyballlover55):

hey so the distance for venus is 108,208,930 snd the distance for jupiter is 778,547,200.

OpenStudy (volleyballlover55):

now what do i do with that

OpenStudy (mayankdevnani):

subtract them first

OpenStudy (mayankdevnani):

do you know why subtract first ?

OpenStudy (volleyballlover55):

ok

OpenStudy (countryboy15):

The distance varies between 670.1 million kilometers at their closest and 1172.7 million kilometers at their farthest. hope this helps also please medal if it did thanks have a great day

OpenStudy (volleyballlover55):

i got 670338270

OpenStudy (volleyballlover55):

ok @CountryBoy15 thanks!

OpenStudy (volleyballlover55):

so the complete answer is .. If a spacecraft was parked on Venus and needed to make a flight to Jupiter, it would need to travel 670,338,270 km.

OpenStudy (volleyballlover55):

@CountryBoy15

OpenStudy (volleyballlover55):

@CountryBoy15

OpenStudy (volleyballlover55):

@CountryBoy15

OpenStudy (countryboy15):

okay what ???

OpenStudy (volleyballlover55):

so the complete answer is .. If a spacecraft was parked on Venus and needed to make a flight to Jupiter, it would need to travel 670,338,270 km. @CountryBoy15

OpenStudy (countryboy15):

yes that sounds good

OpenStudy (volleyballlover55):

ok!

OpenStudy (volleyballlover55):

whats 670,338,270 in scientific notation @CountryBoy15

OpenStudy (volleyballlover55):

@CountryBoy15

OpenStudy (countryboy15):

?

OpenStudy (volleyballlover55):

whats 670,338,270 in scientific notation

OpenStudy (countryboy15):

hmmmmmmmmmmm

OpenStudy (countryboy15):

Here are some examples of scientific notation. 10000 = 1 x 104 24327 = 2.4327 x 104 1000 = 1 x 103 7354 = 7.354 x 103 100 = 1 x 102 482 = 4.82 x 102 10 = 1 x 101 89 = 8.9 x 101 (not usually done) 1 = 100 1/10 = 0.1 = 1 x 10-1 0.32 = 3.2 x 10-1 (not usually done) 1/100 = 0.01 = 1 x 10-2 0.053 = 5.3 x 10-2 1/1000 = 0.001 = 1 x 10-3 0.0078 = 7.8 x 10-3 1/10000 = 0.0001 = 1 x 10-4 0.00044 = 4.4 x 10-4 @volleyballlover55

OpenStudy (countryboy15):

source: https://www.chem.tamu.edu/class/fyp/mathrev/mr-scnot.html

OpenStudy (volleyballlover55):

is this one.... 670,338,270= 6.70338270 x 10^8?

OpenStudy (countryboy15):

sure could be one

OpenStudy (volleyballlover55):

is that correct?

OpenStudy (countryboy15):

anything els please let me know ill be geting off in 1hr or 2 hr so let me know soon and have a great night talk to you later feel free to tag me or ask me tomorrow :)

OpenStudy (countryboy15):

ill look 1 second okay

OpenStudy (countryboy15):

@volleyballlover55 yes it is

OpenStudy (volleyballlover55):

ok thanks :)

OpenStudy (volleyballlover55):

i have some more questions...